# QT-TDM: Planning with Transformer Dynamics Model and Autoregressive Q-Learning

### v1

@@ IV-C Planning
We evaluate the proposed QT-TDM model on real-time planning with MPC, where inference speed needs to be taken into consideration. The inference time grows with the planning horizon $H$, the number of planning samples $J$, and the dimensionality of the environment $\mathcal{D}$. While Transformers serve as large, expressive, and robust dynamics models, they are not optimized for fast inference [15]. The per-dimension tokenization and the autoregressive token prediction lead to a slow inference over long horizons. To solve this issue and achieve faster inference, we use a short planning horizon and employ the Q-Transformer model to estimate a terminal Q-value [17] that provides a long-term return beyond the short-term horizon.
During planning with MPC, we sample $J$ action sequences of length $H$ from a time-dependent multivariate diagonal Gaussian distribution initialized by $(\mu^{0},\sigma^{0})$. Then, trajectories are generated using rollouts from the learned dynamics model (TDM), and the total return $\mathcal{F}_{\varGamma}$ of a trajectory $\varGamma$ is computed as follows:
@@
where $Q_{\phi}(\hat{s}_{H},a_{H}^{N})$ is the terminal Q-value of the last action dimension $a_{H}^{N}$. The distribution $\mu$ and $\sigma$ are updated to the top-$k$ trajectories with the highest total returns. After a fixed number of iterations $\mathcal{I}$, the planning procedure terminates and the first action is executed in a trajectory sampled from the final updated distribution. In addition to sampling from the Gaussian distribution, we also sample $J_{QT}$ action sequences from the learned Q-Transformer model. The planning procedure is summarized in Algorithm 2 and shown in Fig. 1
@@ V Experiments
Benchmarks. We evaluate the performance of QT-TDM model on diverse state-based continuous control tasks from two benchmarks: DeepMind Control Suite (DMC) [19] and MetaWorld [20]. From DMC, we choose two high-dimensional locomotion tasks (Walker Walk and Cheetah Run) and a sparse reward task (Reacher Easy). MetaWorld contains 50 different robotic manipulation tasks, and because of time and computational constraints, we choose six tasks with various challenges. All tasks are shown in Fig. 3.
Baselines. Since the Generalist TDM [15] is the first Transformer-based model to perform continuous real-time planning, it serves as an eligible baseline. However, a comparison with it was not possible because its implementation is not publicly accessible. We compare the performance of QT-TDM against its two individual modules (QT and TDM) to serve as an ablation study as well. Q-Transformer [16] is a Transformer-based model-free algorithm that uses an autoregressive Q-Learning. We provide an extensive evaluation of QT on diverse tasks in an online RL scenarios. TDM is a model-based algorithm that performs real-time planning but without the guidance from a terminal value function.
Experimental Setup. We list all the environment details for all the tasks from the two benchmarks in Table I. For a fair comparison between the two model-based algorithms (QT-TDM and TDM), we use the same planning parameters shown in Table II. All the compared models are evaluated after every $10$K environment steps averaging over $10$ episodes.
Computational Resources. For each task, we trained our method and the baselines with 3 different random seeds. We ran our experiments with 6 Nvidia Quadro 6000 GPUs (24GB) using one GPU for one seed. For one DMC task, the total training and evaluation of our method takes on average 2 days while TDM takes 1.5 days. For one MetaWorld task, our method takes on average 4 days while TDM takes 3.5 days. The model-free QT takes 2 and 6 hours for training one DMC task and one MetaWorld task respectively.
TABLE I: Environment details for the two domains: DMC and MetaWorld. We use action repeat of $4$ for DMC tasks except for Walker, where action repeat of $2$ is used.
@@
TABLE II: Planning parameters used for all tasks.
Parameter MPC
Intial parameters $(\mu^{0},\sigma^{0})$ $(0,2)$
Planning horizon $H$ $3$
Num. of samples $J$ $512$
Num. of iterations $\mathcal{I}$ $6$
Num. of top-$k$ traj. $64$
Figure 4: Learning curves. Three tasks from DMC (top row), episode return as performance metric. Six tasks from MetaWorld (middle and bottom rows), success rate (%) as performance metric. Mean over 3 seeds; shaded areas are standard deviations.
V-A Results
@@
Planning Efficiency. The two compared model-based algorithms (QT-TDM and TDM) perform real-time planning with a myopic planning horizon $(H=3)$. However, QT-TDM relies on a learned terminal Q-value to guide the short-horizon planning. In DMC tasks, TDM fails to solve the Walker task, its learning stagnates at approximately $200$ returns after $100$K environment steps for the Cheetah task, and it relatively solves the sparse reward Reacher task at approximately $600$ returns. In contrast, our proposed QT-TDM model successfully solves all tasks, except for the Cheetah task where it struggles a bit achieving approximately $400$ returns. In MetaWorld tasks, while TDM struggles to solve hard tasks such as Hammer, Door Unlock, and Button Press Wall, QT-TDM successfully solves all six tasks. QT-TDM outperforms TDM with only a 1.3×1.3$\times$1.3 × increase in running time (e.g., from 1.5 days to 2 days for DMC tasks). This is more efficient than the over 2×2$\times$2 × increase in running time required when extending the planning horizon $(H\geq 6)$. This demonstrates that our proposed QT-TDM achieves efficient real-time planning in terms of both performance and computational demands.
Planning vs. Policy. The compared model-free Q-Transformer selects actions with a value-based policy by maximizing Q-values over the discretized bins for all action dimensions. The QT model successfully solves all tasks from MetaWorld, but with less sample efficiency than our QT-TDM model. In DMC tasks with high-dimensional action spaces (Walker and Cheetah), QT was extremely sample-inefficient compared to QT-TDM. In the Walker task, QT achieves approximately 150 returns at 100K environment steps and 600 returns at 500K environment steps, compared with our proposed QT-TDM that achieves approximately 600 returns at 100K environment steps and 900 returns at 500K environment steps. It is expected that the model-based algorithm is more sample-efficient than its model-free counterpart [3], [4]. Nevertheless, the results demonstrate that Q-Transformer [16] is a capable model-free algorithm that can perform effectively in both online and offline RL scenarios with sparse and dense rewards.
V-B Implementation
@@
Input sequence $\mathit{L}(N+1)$ tokens $N$ tokens
@@
V-C Complexity Analysis
We compare the complexity of our QT-TDM model against the Generalist TDM model [15] in terms of model size (i.e., number of parameters) and inference speed. Since the implementation of Generalist TDM is not available, we do not use quantitative measures for inference speed such as wall-time or FLOPs. Instead, we measure inference speed based on planning horizon $H$, number of planning samples $J$ and number of tokens per timestep $\mathcal{T}$ (see Table V). Due to the per-dimension tokenization, Generalist TDM requires $(M+N+1)$ tokens per timestep: $M$ state tokens, $N$ action tokens, and one reward token. In contrast, QT-TDM requires only $(1+N)$ tokens per timestep by reducing the state tokens to a single token using a learned linear layer and by not using the reward token. Additionally, our model utilizes an $85$% shorter planning horizon compared to the Generalist TDM model. Consequently, QT-TDM achieves faster inference speed with $92$% fewer parameters than Generalist TDM. The computational demands of handling a high number of samples $J$ can be mitigated by increasing parallelization (using multiple cores) [15].
@@ VI Conclusion and future work
Computational efficiency. Although QT-TDM demonstrates fast and computationally efficient inference, the training process still suffers from computational inefficiencies, a known limitation of Transformer-based methods. The primary cause of this inefficiency is the computationally expensive self-attention layer, whose complexity grows quadratically with the number of tokens. The GFNet model [31] replaces the self-attenstion layer in the vision Transformer with a proposed global filter layer. The global filter layer learns the spatial features in the frequency domain by applying a 2D Discrete Fourier Transform (DFT) with log-linear complexity. A recently proposed model, TSLANet [32], replaces the self attention layer with a lightweight Adaptive Spectral Block (ASB) that employs 1D DFT to learn representations in the frequency domain. TSLANet outperforms Transformer-based models in time-series analysis while utilizing significantly fewer parameters and fewer FLOPs. We plan to test learning representations in the frequency domain using DFT in the reinforcement learning domain, anticipating that it will deliver comparable performance to that observed in vision and time series domains.
@@
Acknowledgement. The authors gratefully acknowledge support from the German Research Foundation DFG under project CML (TRR 169). Mostafa Kotb is funded by a scholarship from the Ministry of Higher Education of the Arab Republic of Egypt.

### v2

@@ IV-C Planning
We evaluate the proposed QT-TDM model on real-time planning with MPC, where inference speed needs to be taken into consideration. The inference time grows with the planning horizon $H$, the number of planning samples $J$, and the dimensionality of the environment $\mathcal{D}$. While Transformers serve as large, expressive, and robust dynamics models, they are not optimized for fast inference [12]. The per-dimension tokenization and the autoregressive token prediction lead to a slow inference over long horizons. To solve this issue and achieve faster inference, we use a short planning horizon and employ the Q-Transformer model to estimate a terminal Q-value [19] that provides a long-term return beyond the short-term horizon. During planning with MPC, we sample $J$ action sequences of length $H$ from a time-dependent multivariate diagonal Gaussian distribution initialized by $(\mu^{0},\sigma^{0})_{t:t+H}$. Then, trajectories are generated using rollouts from the learned dynamics model (TDM), and the total return $\mathcal{F}_{\varGamma}$ of a trajectory $\varGamma$ is computed as follows:
@@
where $Q_{\phi}(\hat{s}_{H},a_{H}^{N})$ is the terminal Q-value of the last action dimension $a_{H}^{N}$. The distribution $\mu^{n}$ and $\sigma^{n}$ at iteration $n$ are updated to the top-$k$ trajectories with the highest total returns $\mathcal{F}_{\varGamma}^{\star}$ as follows:
$\mu^{n}=\frac{\sum_{i=1}^{k}\Omega_{i}\varGamma_{i}^{\star}}{\sum_{i=1}^{k}\Omega_{i}},\quad\sigma^{n}=\sqrt{\frac{\sum_{i=1}^{k}\Omega_{i}\left(\varGamma_{i}^{\star}-\mu^{n}\right)^{2}}{\sum_{i=1}^{k}\Omega_{i}}},$ (9)
where $\Omega_{i}=e^{\tau\left(\mathcal{F}_{\varGamma,i}^{\star}\right)}$, $\tau$ is a temperature parameter controlling the sharpness of the weighting and $\varGamma_{i}^{\star}$ is the $i$th top-$k$ trajectory. After a fixed number of iterations $\mathcal{I}$, the planning procedure terminates and a trajectory is sampled from the final updated distribution. The first action is only executed as we plan at each decision step $t$. In addition to sampling from the Gaussian distribution, we also sample $J_{QT}$ action sequences from the learned Q-Transformer model. The planning procedure is summarized in Algorithm 2 and shown in Fig. 1
@@ V Experiments
V-A Description and Details
V-A1 Benchmarks
We evaluate the performance of QT-TDM model on diverse state-based continuous control tasks from two benchmarks: DeepMind Control Suite (DMC) [20] and MetaWorld [21]. From DMC, we choose two high-dimensional locomotion tasks (Walker Walk and Cheetah Run) and a sparse reward task (Reacher Easy). MetaWorld contains 50 different robotic manipulation tasks, and because of time and computational constraints, we choose six tasks with various challenges. All tasks are shown in Fig. 3.
V-A2 Baselines
Since the Generalist TDM [12] is the first Transformer-based model to perform continuous real-time planning, it serves as an eligible baseline. However, a comparison with it was not possible because its implementation is not publicly accessible. We compare the performance of QT-TDM against PlaNet [22], DreamerV3 [33] and its two individual modules (QT and TDM) to serve as an ablation study as well. Both, PlaNet and DreamerV3 are model-based algorithms that use Recurrent State-Space Model (RSSM) as dynamics model. While PlaNet performs real-time planning with MPC, DreamerV3 performs background planning. Q-Transformer [18] is a Transformer-based model-free algorithm that uses an autoregressive Q-Learning. We provide an extensive evaluation of QT on diverse tasks in an online RL scenarios. TDM is a model-based algorithm that performs real-time planning but without the guidance from a terminal value function.
V-A3 Experimental Setup
We list all the environment details for the tasks from the two benchmarks in Table I . For a fair comparison between the two Transformer-based model-based algorithms (QT-TDM and TDM), we use the same planning parameters shown in Table II. For the Recurrent-based model-based algorithm (PlaNet), we use its default planning parameters. All the compared models are evaluated after every $10$K environment steps averaging over $10$ episodes, except for DreamerV3, for which we use the final performance after convergence that we obtained from [34].
V-A4 Computational Resources
For each task, we trained our method and the baselines with 3 different random seeds. We ran our experiments with 6 Nvidia Quadro 6000 GPUs (24GB) using one GPU for one seed. For one DMC task, the total training and evaluation of our method takes on average 2 days while TDM takes 1.5 days. For one MetaWorld task, our method takes on average 4 days while TDM takes 3.5 days. The model-free QT takes 2 and 6 hours for training one DMC task and one MetaWorld task respectively.
TABLE I: Environment details used across all methods for the two domains. We use action repeat of $4$ for DMC tasks except for Walker, where action repeat of $2$ is used.
@@ V-A4 Computational Resources
TABLE II: MPC Planning parameters used for all tasks.
Parameter QT-TDM (ours) / TDM PlaNet [22]
Initial parameters $(\mu^{0},\sigma^{0})$ $(0,2)$ $(0,1)$
Planning horizon $H$ $3$ $12$
Num. of samples $J$ $512$ $1000$
Num. of iterations $\mathcal{I}$ $6$ $10$
Num. of top-$k$ trajectories $64$ $100$
Figure 4: Learning curves. Three tasks from DMC (top row), episode return as performance metric. Six tasks from MetaWorld (middle and bottom rows), success rate (%) as performance metric. Mean over 3 seeds; shaded areas are standard deviations. For DreamerV3, we report the final performance from [34].
V-B Results
@@
V-B1 Planning Efficiency
The two compared Transformer-based model-based algorithms (QT-TDM and TDM) perform real-time planning with a myopic planning horizon $(H=3)$. However, QT-TDM relies on a learned terminal Q-value to guide the short-horizon planning. In DMC tasks, TDM fails to solve the Walker task, its learning stagnates at approximately $200$ returns after $100$K environment steps for the Cheetah task, and it relatively solves the sparse reward Reacher task at approximately $600$ returns. In contrast, our proposed QT-TDM model successfully solves all tasks, except for the Cheetah task where it struggles a bit achieving approximately $400$ returns. We achieved improved results with planning horizons $H=5$ and $H=9$ as shown in Fig. 5 but with the cost of higher inference time. In MetaWorld tasks, while TDM struggles to solve hard tasks such as Hammer, Door Unlock, and Button Press Wall, QT-TDM successfully solves all six tasks. QT-TDM outperforms TDM with only a 1.3×1.3$\times$1.3 × increase in running time (e.g., from 1.5 days to 2 days for DMC tasks). This is more efficient than the over 2×2$\times$2 × increase in running time required when extending the planning horizon $(H\geq 6)$. This demonstrates that our proposed QT-TDM achieves efficient real-time planning in terms of both performance and computational demands.
V-B2 Transformer vs. Recurrent
Comparing our Transformer-based model against two Recurrent-based models highlights the superiority of TDMs in modeling dynamics. QT-TDM consistently outperforms PlaNet across all tasks, even though PlaNet utilizes a longer planning horizon (see table II). When compared to the state-of-the-art DreamerV3 which performs background planning, QT-TDM surpasses it in all MetaWorld tasks, while DreamerV3 achieves better performance in two DMC tasks (in Reacher Easy and Cheetah Run).
V-B3 Planning vs. Policy
The compared model-free Q-Transformer selects actions with a value-based policy by maximizing Q-values over the discretized bins for all action dimensions. The QT model successfully solves all tasks from MetaWorld, but with less sample efficiency than our QT-TDM model. In DMC tasks with high-dimensional action spaces (Walker and Cheetah), QT was extremely sample-inefficient compared to QT-TDM. In the Walker task, QT achieves approximately 150 returns at 100K environment steps and 600 returns at 500K environment steps, compared with our proposed QT-TDM that achieves approximately 600 returns at 100K environment steps and 900 returns at 500K environment steps. It is expected that the model-based algorithm is more sample-efficient than its model-free counterpart [3], [4]. Nevertheless, the results demonstrate that QT [18] is a capable model-free algorithm that can perform effectively in both online and offline RL scenarios with sparse and dense rewards.
V-C Implementation
@@
Figure 5: QT-TDM with different planning horizon $(H)$ on Cheetah Run task.
@@
Input sequence $\mathit{L}\times(N+1)$ tokens $N$ tokens
@@
V-D Complexity Analysis
We compare the complexity of our QT-TDM model against the Generalist TDM model [12] in terms of model size (i.e., number of parameters) and inference speed. Since the implementation of Generalist TDM is not available, we do not use quantitative measures for inference speed such as wall-time or FLOPs. Instead, we measure inference speed based on planning horizon $H$, terminal value $Q^{H}$, number of planning samples $J$ and number of tokens per timestep $\mathcal{T}$ (see Table V). Due to the per-dimension tokenization, Generalist TDM requires $(M+N+1)$ tokens per timestep: $M$ state tokens, $N$ action tokens, and one reward token. In contrast, QT-TDM requires only $(1+N)$ tokens per timestep by reducing the state tokens to a single token using a learned linear layer and by not using the reward token. Additionally, our model utilizes an $85$% shorter planning horizon compared to the Generalist TDM model. However, our model leverages a terminal Q-value $N$ times, with one value for each action dimension. Despite the additional steps, the total planning steps required by our model ($3\!+\!N$ steps) remain fewer than those required by the Generalist TDM (at least $20$ steps). Consequently, QT-TDM achieves faster inference speed with $92$% fewer parameters than Generalist TDM. The computational demands of handling a high number of samples $J$ can be mitigated by increasing parallelization (using multiple cores) [12].
@@ VI Conclusion and future work
Acknowledgment
The authors would like to thank Ozan Özdemir for his suggestions during the revision phase.